Diffusion Morning Report 04/03/12

Ted Berg tberg at stanford.edu
Tue Apr 3 11:35:17 PDT 2012


Good morning All,
     Looking OK this morning. New items:LTO is down for pull and clean.
     Leftovers: Thermco TLC issues,SiGe pump out for repair,POCl3 
andBBr3 sensor issues,Epi chamber 2 install-ongoing,TEOS-mfs issues,Thermco
Nitride leak. No other major issues at this time.


					
	Date: 	4/3/12 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*yellow* 	*tgo sensor issue-ran BBr3 run-observed* 	** 
** 	* 2*
Tylan 6-POCl3 	*yellow* 	*tgo sensor issue tube reinstalled-ran POCl3 
run* 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*mfs error* 	*14.39/30* 	** 	**
Tylan 8-SiGe 	*down* 	*broken tube-bad pump-out for repair* 
*3.4/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*15.78/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*14.22/30* 	** 	**
Tylan 11-LTO 	*down* 	*pull and clean* 	*0./30* 	** 	* 1*
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*yellow* 	*no tlc* 	** 	** 	* 3*
Thermco 2 	*yellow* 	*no tlc* 	** 	** 	**
Thermco 3 	*yellow* 	*testing TLC* 	** 	** 	**
Thermco4 	*up* 	** 	** 	** 	**
Thermco Poly1 	*yellow* 	*no B2H6-reg ordered* 	** 	*
* 	**
Thermco Poly2 	*yellow* 	*no B2H6-reg ordered* 	** 	** 	**
Thermco Nitride1 	*yellow* 	*leak waiting for o-rings* 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*down* 	*starting up,scrubber prob on Chamber 2* 	** 	** 	**
TEL Spa 	*yellow* 	*needs test* 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
SEM 4160 	*up* 	** 	** 	** 	**
ZYGO 	*up* 	** 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	*on* 	*crossover ok* 	*changeover working well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20120403/004a2f98/attachment.html>


More information about the diffusion mailing list